# Multi: Multimodal Understanding Leaderboard with Text and Images

### v2

@@ Appendix B Models
Creator Model # Paras Form Modality Lang Version
FDU MOSS Sun et al. (2023b) 16B Weight T zh, en moss-moon-003-sft
SJTU&AISpeech DFM-2.0 Chen et al. (2022) 70B Weight T zh, en dfm-2.0-70b-preview
LinkSoul-AI Chinese-LLaVA LinkSoul-AI (2023) 7B Weight One zh, en Chinese-LLaVA-Cllama2
THU VisualGLM Du et al. (2022) 6B Weight SI zh, en visualglm-6b
ModelBest VisCPM Hu et al. (2023) 10B Weight SI zh, en VisCPM-Chat
Alibaba Qwen-VL Bai et al. (2023a) 7B Weight MI zh, en Qwen-VL-Chat
OpenGVLab InternVL Chen et al. (2023) 19B Weight One zh, en InternVL-Chat-Chinese-V1.1
01-ai Yi-VL 01.ai (2023) 34B Weight One zh, en Yi-34B-Chat
Google Gemini Team (2023) - API T ML gemini-pro
Google Gemini Vision Team (2023) - API MI ML gemini-pro-vision
OpenAI ChatGPT OpenAI (2022) - API T ML gpt-3.5-turbo-1106
OpenAI GPT-4 OpenAI (2023a) - API T ML gpt-4-1106-preview
OpenAI GPT-4V OpenAI (2023b) - API MI ML gpt-4-vision-preview
Table 10: The list of models evaluated on Multi. We report Modality as how many images can the model take in one turn. Note that those MLLMs commonly support multiple-image input with chatting in several turns. W: accessible through weight. T: pure text LLM, One: only one image in the beginning, SI: single image in each turn, MI: multiple images in one turn. The underline means the model must have an image as input. ML: Multi-lingual.
@@ Appendix D Data Selection Algorithm
$L_{q}=$ $\left(a\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{characters in question}})\\ \mathcal{H}(L_{q,\#\text{characters in answer}})\\ \mathcal{H}(L_{q,\#\text{characters in analysis}})\end{bmatrix}\right.+$
$\left.b\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{images in question}})\\ \mathcal{H}(L_{q,\#\text{images in answer}})\\ \mathcal{H}(L_{q,\#\text{images in analysis}})\end{bmatrix}\right)^{\top}\begin{bmatrix}1.0\\ 0.1\\ 0.5\end{bmatrix}$
where $q=1,b=1$ are customized weights.
In the formula above, we use a harmonic mean function $\mathcal{H}$ to normalize content length $L_{q,i}$ of each target value $i$ within each knowledge piece $k$ [^6] Note that for those questions without knowledge information, we simply use a ”null” string as a keyword.
@@
$Pr[\text{pick up }q]=\begin{cases}1\text{ , for }q\text{ s.t. }L_{q,k}[0]\\ p\text{ , if }q=1\text{ , for }q\text{ of }L_{q,k}[1:m]\\ \quad\quad\quad\quad\quad\quad\quad\text{ or }L_{q,k}[-m:]\\ p\frac{N_{k}-2m}{\#\text{questions of }k}\text{ , otherwise}\end{cases}$
@@ E.3 Data Post-process
Appendix F More Examples
@@
Figure 9: More example of Multi.

### v3

@@ Appendix B Models and Experiment Cost
Creator Model Date # Paras Form Modality Version
FDU MOSS Sun et al. (2023) 2023-04-19 16B Weight T moss-moon-003-sft
AISpeech DFM-2.0 Chen et al. (2022) 2024-01-16 70B Weight T dfm-2.0-70b-preview
Alibaba Qwen2-7B qwe (2024) 2024-06-04 7B Weight T Qwen2-7B-Instruct
PJLab InternLM2.5-7B Cai et al. (2024) 2024-06-27 7B Weight T internlm2_5-7b-chat
[…]
OpenBMB VisCPM Hu et al. (2023) 2023-06-28 10B Weight SI VisCPM-Chat
LinkSoul-AI Chinese-LLaVA LinkSoul-AI (2023) 2023-06-30 7B Weight One Chinese-LLaVA-Cllama2
Alibaba Qwen-VL Bai et al. (2023a) 2023-09-25 7B Weight MI Qwen-VL-Chat
01-ai Yi-VL-34B 01.ai (2023) 2023-11-22 34B Weight One Yi-VL-34B
OpenGVLab InternVL-1.1 Chen et al. (2024) 2024-01-24 19B Weight One InternVL-Chat-V1-1
OpenGVLab InternVL2-8B Zhe et al. (2024) 2024-07-03 8B Weight MI InternVL2-8B
OpenBMB MiniCPM-V-2.6 Yao et al. (2024) 2024-08-05 8B Weight MI MiniCPM-V-2_6
Alibaba Qwen2-VL-7B Wang et al. (2024) 2024-08-28 7B Weight MI Qwen2-VL-7B-Instruct
Alibaba Qwen2-VL-72B Wang et al. (2024) 2024-08-28 72B Weight MI Qwen2-VL-72B-Instruct
OpenAI ChatGPT OpenAI (2022) 2023-11-06 - API T gpt-3.5-turbo-1106
OpenAI GPT-4 OpenAI (2023a) 2023-11-06 - API T gpt-4-1106-preview
Google Gemini Pro Team (2023) 2023-12-07 - API T gemini-1.0-pro-001
[…]
Google Gemini Vision Pro Team (2023) 2023-12-07 - API MI gemini-1.0-pro-vision
OpenAI GPT-4o-mini OpenAI (2024a) 2024-07-18 - API MI gpt-4o-mini-2024-07-18
OpenAI GPT-4o (0806) OpenAI (2024b) 2024-08-06 - API MI gpt-4o-2024-08-06
Google Gemini 1.5 Flash Team (2023) 2024-08-27 - API MI gemini-1.5-flash-exp-0827
Anthropic Claude 3.5 Sonnet Anthropic $-2$ API MI claude-3-5-sonnet-20241022
OpenAI GPT-4o (1120) OpenAI (2024b) 2024-11-20 - API MI gpt-4o-2024-11-20
Google Gemini 1.5 Pro Team (2023) 2024-11-21 - API MI gemini-exp-1121
Table 10: The list of models evaluated on Multi. We report Modality as how many images can the model take in one turn. Note that those MLLMs commonly support multiple-image input with chatting in several turns. T: pure text LLM, One: only one image in the beginning, SI: single image in each turn, MI: multiple images in one turn. The underline means the model must have an image as input.
A standard evaluation using GPT-4o with image inputs (without CoT reasoning) requires approximately 4.5 million prompt tokens and 150,000 completion tokens, amounting to a total cost of around \$25. Running on a single A100(80G) GPU, models with 7–8B parameters complete the benchmark in 1–2 hours, while a 70B+ model takes approximately 8 hours on 2-4 A100 GPUs. CoT experiments demand significantly more resources due to the larger token output. To facilitate more practical and faster usage, we recommend researchers prioritize testing on MULTI-Elite, which requires only around \$1 and a few minutes to complete.
@@ Appendix D Data Selection Algorithm
$L_{q}=\left(a\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{characters in question}})\\ \mathcal{H}(L_{q,\#\text{characters in answer}})\\ \mathcal{H}(L_{q,\#\text{characters in analysis}})\end{bmatrix}\right.+\left.b\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{images in question}})\\ \mathcal{H}(L_{q,\#\text{images in answer}})\\ \mathcal{H}(L_{q,\#\text{images in analysis}})\end{bmatrix}\right)^{\top}\begin{bmatrix}1.0\\ 0.1\\ 0.5\end{bmatrix}$
where $a=1,b=1$ are customized weights.
In the formula above, we use a harmonic mean function $\mathcal{H}$ to normalize content length $L_{q,i}$ of each target value $i$ within each knowledge piece $k$ [^5] Note that for those questions without knowledge information, we simply use an empty string as a keyword.
@@
$Pr[\text{pick up }q]=\begin{cases}1&\text{ , for }q\text{ s.t. }L_{q,k}[0]\\ p&\text{ , if }q=1\text{ , for }q\text{ of }L_{q,k}[1:m]\text{ or }L_{q,k}[-m:]\\ p\frac{N_{k}-2m}{\#\text{questions of }k}&\text{ , otherwise}\end{cases}$
@@ E.3 Data Post-process
Figure 14: More example of Multi.
@@
Appendix F Case study
In this section, we provide 5 case studies alone with annotations by human experts and error analysis for several selected models.
Figure 15: Error analysis of a SI-SA question.
Figure 16: Error analysis of a SI-SA question.
Figure 17: Error analysis of a MI-SA question.
Figure 18: Error analysis of a SI-OP question.
Figure 19: Error analysis of a NI-MA question.